# Large Language Models for Information Retrieval: A Survey

### v2

@@ 5.3 Utilizing LLMs for Training Data Augmentation
Furthermore, in the realm of reranking, researchers have explored the integration of LLMs for training data augmentation [133, 134, 135, 130]. For example, ExaRanker [133] generates explanations for retrieval datasets using GPT-3.5, subsequently trains a seq2seq ranking model to generate relevance labels along with corresponding explanations for given query-document pairs. InPars-Light [134] is proposed as a cost-effective method to synthesize queries for documents by prompting LLMs. Contrary to InPars-Light [134], a new dataset ChatGPT-RetrievalQA [135] is constructed by generating synthetic documents based on LLMs in response to user queries. Besides, there have been attempts to distill the document ranking capability of ChatGPT into a specialized model [130]. In this approach, ChatGPT is first instructed to directly generate a ranking list of the documents. Then, the generated document list is used as the target to train a student model (i.e., DeBERTa-v3-base) using various ranking losses (e.g., RankNet [125]).
@@ 6 Reader
TABLE V: The comparison of existing methods that have a passive reader module. REALM and RAG do not use LLMs, but their frameworks have been widely applied in many following approaches.
@@
RETA-LLM [143] LLaMA & GLM & GPT Input layer In the beginning Prompting
@@
A naive strategy for implementing this function is to heuristically provide LLMs with documents relevant to the user queries or the previously generated texts to support following generation. However, this passive approach limits LLMs to merely collecting documents from IR systems without active engagement. An alternative solution is to train LLMs to interact proactively with search engines. For example, LLMs can formulate their own queries instead of relying solely on user queries or generated texts for references. According to the way LLMs utilize IR systems in the reader module, we can categorize them into passive reader and active reader. Each approach has its advantages and challenges for implementing LLM-powered answer generation in IR systems.
@@ 6.1 Passive Reader
To generate answers for users, a straightforward strategy is to supply the retrieved documents according to the queries or previously generated texts from IR systems as inputs to LLMs for creating passages [143, 137, 139, 140, 141, 142, 145, 138, 144, 23, 146]. By this means, these approaches use the LLMs and IR systems in a separate manner, with LLMs functioning as passive recipients of documents from the IR systems. The strategies for utilizing LLMs within IR systems’ reader modules can be categorized into the following three groups according to the frequency of retrieving documents for LLMs.
@@ 6.1.1 Once-Retrieval Reader
To obtain useful references for LLMs to generate responses for user queries, an intuitive way is to retrieve the top documents based on the queries themselves in the beginning. For example, REALM [137] adopts this strategy by directly attending the document contents to the original queries to predict the final answers based on masked language modeling. RAG [138] follows this strategy but applies the generative language modeling paradigm. However, these two approaches only use language models with limited parameters, such as BERT and BART. Recent approaches such as REPLUG [139] and Atlas [140] have improved them by leveraging LLMs such as GPTs and T5s for response generation. To yield better answer generation performances, these models usually fine-tune LLMs on QA tasks. However, due to the limited computing resources, many methods [141, 142] choose to prompt LLMs for generation as they could use larger LMs in this way. To better facilitate fixed LLMs to perform the reading tasks, RETA-LLM [143] breaks down the whole complex generation task into several simple modules in the reader pipeline. These modules include a query rewriting module for refining query intents, a passage extraction module for aligning reference lengths with LLM limitations, and a fact verification module for confirming the absence of fabricated information in the generated answers.
@@ 6.1.2 Periodic-Retrieval Reader
However, while generating long conclusive answers, it is shown [23, 144] that only using the references retrieved by the original user intents as in once-retrieval readers may be inadequate. For example, when providing a passage about “Barack Obama”, language models may need additional knowledge about his university, which may not be included in the results of simply searching the initial query. In conclusion, language models may need extra references to support the following generation during the generating process, where multiple retrieval processes may be required. To address this, solutions such as RETRO [23] and RALM [144] have emerged, emphasizing the periodic collection of documents based on both the original queries and the concurrently generated texts (triggering retrieval every $n$ generated tokens). In this manner, when generating the text about the university career of Barack Obama, the LLM can receive additional documents as supplementary materials. This need for additional references highlights the necessity for multiple retrieval iterations to ensure robustness in subsequent answer generation. Notably, RETRO [23] introduces a novel approach incorporating cross-attention between the generating texts and the references within the Transformer attention calculation, as opposed to directly embedding references into the input texts of LLMs. Since it involves additional cross-attention modules in the Transformer’s structure, RETRO trains this model from scratch. However, these two approaches mainly rely on the successive $n$ tokens to separate generation and retrieve documents, which may not be semantically continuous and may cause the collected references noisy and useless. To solve this problem, some approaches such as IRCoT [145] also explore retrieving documents for every generated sentence, which is a more complete semantic structure.
@@ 6.1.3 Aperiodic-Retrieval Reader
In the above strategy, the retrieval systems supply documents to LLMs in a periodic manner. However, retrieving documents in a mandatory frequency may mismatch the retrieval timing and can be costly. Recently, FLARE [146] has addressed this problem by automatically determining the timing of retrieval according to the probability of generating texts. Since the probability can serve as an indicator of LLMs’ confidence during text generation [147, 148], a low probability for a generated term could suggest that LLMs require additional knowledge. Specifically, when the probability of a term falls below a predefined threshold, FLARE employs IR systems to retrieve references in accordance with the ongoing generated sentences, while removing these low-probability terms. FLARE adopts this strategy of prompting LLMs for answer generation solely based on the probabilities of generating terms, avoiding the need for fine-tuning while still maintaining effectiveness.
We summarize these passive reader approaches in Table V, considering various aspects such as the backbone language models, the insertion point for retrieved references, the timing of using retrieval models, and the tuning strategy employed for LLMs.
@@ 6.2 Active Reader
To allow LLMs to actively use search engines, Self-Ask [149] and DSP [150] try to employ few-shot prompts for LLMs, triggering them to search queries when they believe it is required. For example, in a scenario where the query is “When was the existing tallest wooden lattice tower built?”, these prompted LLMs can decide to search a query “What is the existing tallest wooden lattice tower” to gather necessary references as they find the query cannot be directly answered. Once acquired information about the tower, they can iteratively query IR systems for more details until they determine to generate the final answers instead of asking questions. Notably, these methods involve IR systems to construct a single reasoning chain for LLMs. MRC [151] further improves these methods by prompting LLMs to explore multiple reasoning chains and subsequently combining all generated answers together using LLMs.
Instead of prompting LLMs, WebGPT [24] takes an alternate approach by training GPT-3 models to use search engines automatically. This is achieved through the application of a reinforcement learning framework, within which a simulated environment is constructed for GPT-3 models. Specifically, the WebGPT model employs special tokens to execute actions such as querying, scrolling through rankings, and quoting references on search engines. This innovative approach allows the GPT-3 model to use search engines for text generation, enhancing the reliability and real-time capability of the generated texts. Some following models [152] have extended this paradigm to the domain of Chinese question answering.
@@ 6.3 Limitations
Several IR systems applying the retrieval-augmented generation strategy, such as New Bing and Langchain, have already entered commercial use. However, there are also some challenges in this novel retrieval-augmented content generation system. These include challenges such as effective query reformulation, optimal retrieval frequency, correct document comprehension, accurate passage extraction, and effective content summarization. It is crucial to address these challenges in order to effectively realize the potential of LLMs in this paradigm.
@@ 7.1 Query Rewriter
$\bullet$ Rewriting query according to ranking performance. A typical paradigm of prompting-based methods is providing LLMs with several ground-truth rewriting cases (optional) and the task description of query rewriting. Despite LLMs being capable of identifying potential user intents of the query [153], they lack awareness of the resulting retrieval quality of the rewritten query. The absence of this connection can result in rewritten queries that seem correct yet produce unsatisfactory ranking results. Although some existing studies have used reinforcement learning to adjust the query rewriting process according to generation results [94], a substantial realm of research remains unexplored concerning the integration of ranking results.
@@
$\bullet$ Achieving personalized query rewriting. LLMs offer valuable contributions to personalized search through their capacity in analyzing user-specific data. In terms of query rewriting, with the excellent language comprehension ability of LLMs, it is possible to leverage them to build user profiles based on users’ search histories (e.g., issued queries, click-through behaviors, and dwell time). This empowers the achievement of personalized query rewriting for enhanced information retrieval and finally benefits personalized search or personalized recommendation.

### v3

@@ 5.3 Utilizing LLMs for Training Data Augmentation
Furthermore, in the realm of reranking, researchers have explored the integration of LLMs for training data augmentation [157, 158, 159, 160, 161, 162]. For example, ExaRanker [157] generates explanations for retrieval datasets using GPT-3.5, and subsequently trains a seq2seq ranking model to generate relevance labels along with corresponding explanations for given query-document pairs. InPars-Light [158] is proposed as a cost-effective method to synthesize queries for documents by prompting LLMs. Contrary to InPars-Light [158], a new dataset ChatGPT-RetrievalQA [159] is constructed by generating synthetic documents based on LLMs in response to user queries.
Recently, many studies [160, 161, 162] have also attempted to distill the document ranking capability of LLMs into a specialized model. RankVicuna [160] proposes to use the ranking list of $\text{RankGPT}_{3.5}$ [152] as the gold list to train a 7B parameter Vicuna model. RankZephyr [161] introduces a two-stage training strategy for distillation: initially applying the RankVicuna recipe to train $\text{Zephyr}{\gamma}$ in the first stage, and then further finetuning it in the second stage with the ranking results from $\text{RankGPT}_{4}$. These two studies not only demonstrate competitive results but also alleviate the issue of ranking results non-reproducibility of black-box LLMs. Besides, researchers [162] have also tried to distill the ranking ability of a pairwise ranker, which is computationally demanding, into a simpler but more efficient pointwise ranker.
@@ 6 Reader
TABLE VII: The comparison of existing representative methods that have a passive reader module. REALM and RAG do not use LLMs, but their frameworks have been widely applied in many following approaches.
@@
Chain-of-Note [172] LLaMA Input layer In the beginning Fine-tuning
@@
ITERGEN [174] GPT Input layer During generation (every answer) Prompting
@@
Self-RAG [177] LLaMA Input layer During generation (aperiodic) Fine-tuning
@@
A naive strategy for implementing this function is to heuristically provide LLMs with documents relevant to the user queries or the previously generated texts to support the following generation. However, this passive approach limits LLMs to merely collecting documents from IR systems without active engagement. An alternative solution is to train LLMs to interact proactively with search engines. For example, LLMs can formulate their own queries instead of relying solely on user queries or generated texts for references. According to the way LLMs utilize IR systems in the reader module, we can categorize them into passive readers and active readers. Each approach has its advantages and challenges for implementing LLM-powered answer generation in IR systems. Furthermore, since the documents provided by upstream IR systems are sometimes too long to directly feed as input for LLMs, some compression modules are proposed to extractively or abstractively compress the retrieved contexts for LLMs to understand and generate answers for queries. We will present these reader and compressor modules in the following parts and briefly introduce the existing analysis work on retrieval-augmented generation strategy and their applications.
@@ 6.1 Passive Reader
To generate answers for users, a straightforward strategy is to supply the retrieved documents according to the queries or previously generated texts from IR systems as inputs to LLMs for creating passages [178, 166, 168, 169, 170, 171, 179, 180, 175, 167, 173, 23, 176]. By this means, these approaches use the LLMs and IR systems separately, with LLMs functioning as passive recipients of documents from the IR systems. The strategies for utilizing LLMs within IR systems’ reader modules can be categorized into the following three groups according to the frequency of retrieving documents for LLMs.
@@ 6.1.1 Once-Retrieval Reader
To obtain useful references for LLMs to generate responses for user queries, an intuitive way is to retrieve the top documents based on the queries themselves in the beginning. For example, REALM [166] adopts this strategy by directly attending the document contents to the original queries to predict the final answers based on masked language modeling. RAG [167] follows this strategy but applies the generative language modeling paradigm. However, these two approaches only use language models with limited parameters, such as BERT and BART. Recent approaches such as REPLUG [168] and Atlas [169] have improved them by leveraging LLMs such as GPTs, T5s, and LLaMAs for response generation. To yield better answer generation performances, these models usually fine-tune LLMs on QA tasks. However, due to the limited computing resources, many methods [170, 171, 179] choose to prompt LLMs for generation as they could use larger LMs in this way. Furthermore, to improve the quality of the generated answers, several approaches [181, 172] also try to train or prompt the LLMs to generate contexts such as citations or notes in addition to the answers to force LLMs to understand and assess the relevance of retrieved passages to the user queries. Some approaches [180] evaluate the importance of each retrieved reference using policy gradients to indicate which reference is more useful for generating. Besides, researchers explore instruction tuning LLMs such LLaMAs to improve their abilities to generate conclusive passages relying on retrieved knowledge [182, 183].
@@ 6.1.2 Periodic-Retrieval Reader
However, while generating long conclusive answers, it is shown [23, 173] that only using the references retrieved by the original user intents as in once-retrieval readers may be inadequate. For example, when providing a passage about “Barack Obama”, language models may need additional knowledge about his university, which may not be included in the results of simply searching the initial query. In conclusion, language models may need extra references to support the following generation during the generating process, where multiple retrieval processes may be required. To address this, solutions such as RETRO [23] and RALM [173] have emerged, emphasizing the periodic collection of documents based on both the original queries and the concurrently generated texts (triggering a retrieval every $n$ generated tokens). In this manner, when generating the text about the university career of Barack Obama, the LLM can receive additional documents as supplementary materials. This need for additional references highlights the necessity for multiple retrieval iterations to ensure robustness in subsequent answer generation. Notably, RETRO [23] introduces a novel approach incorporating cross-attention between the generating texts and the references within the Transformer attention calculation, as opposed to directly embedding references into the input texts of LLMs. Since it involves additional cross-attention modules in the Transformer’s structure, RETRO trains this model from scratch. However, these two approaches mainly rely on the successive $n$ tokens to separate generation and retrieve documents, which may not be semantically continuous and may cause the collected references noisy and useless. To solve this problem, some approaches such as IRCoT [175] also explore retrieving documents for every generated sentence, which is a more complete semantic structure. Furthermore, researchers find that the whole generated passages can be considered as conclusive contexts for current queries and can be used to find more relevant knowledge to generate more thorough answers. Consequently, many recent approaches [184, 174, 185] have also tried to extend this periodic-retrieval paradigm to iteratively using the whole generated passages to retrieve references to re-generate the answers, until the iterations reach a pre-defined limitation. Particularly, these methods can be regarded as special periodic-retrieval readers that retrieve passages when every answer is (re)-generated. Since the LLMs can receive more comprehensive and relevant references with the iterations increase, these methods that combine retrieval-augmented-generation and generation-augmented-retrieval strategies can generate more accurate answers but consume more computation costs.
@@ 6.1.3 Aperiodic-Retrieval Reader
In the above strategy, the retrieval systems supply documents to LLMs in a periodic manner. However, retrieving documents in a mandatory frequency may mismatch the retrieval timing and can be costly. Recently, FLARE [176] has addressed this problem by automatically determining the timing of retrieval according to the probability of generating texts. Since the probability can serve as an indicator of LLMs’ confidence during text generation [186, 187], a low probability for a generated term could suggest that LLMs require additional knowledge. Specifically, when the probability of a term falls below a predefined threshold, FLARE employs IR systems to retrieve references in accordance with the ongoing generated sentences, while removing these low-probability terms. FLARE adopts this strategy of prompting LLMs for answer generation solely based on the probabilities of generating terms, avoiding the need for fine-tuning while still maintaining effectiveness. Besides, self-RAG [177] tends to solve this problem by training LLMs such as LlaMA to generate specific tokens when they need additional knowledge to support following generations. Another critical model is introduced to judge whether the retrieved references are beneficial for generating.
We summarize representative passive reader approaches in Table VII, considering various aspects such as the backbone language models, the insertion point for retrieved references, the timing of using retrieval models, and the tuning strategy employed for LLMs.
@@ 6.2 Active Reader
To allow LLMs to actively use search engines, Self-Ask [188] and DSP [189] try to employ few-shot prompts for LLMs, triggering them to search queries when they believe it is required. For example, in a scenario where the query is “When was the existing tallest wooden lattice tower built?”, these prompted LLMs can decide to search a query “What is the existing tallest wooden lattice tower” to gather necessary references as they find the query cannot be directly answered. Once acquired information about the tower, they can iteratively query IR systems for more details until they determine to generate the final answers instead of asking questions. Notably, these methods involve IR systems to construct a single reasoning chain for LLMs. MRC [190] further improves these methods by prompting LLMs to explore multiple reasoning chains and subsequently combining all generated answers using LLMs.
6.3 Compressor
Existing LLMs, especially open-sourced ones, such as LLaMA and Flan-T5, have limited input lengths (usually 4,096 or 8,192 tokens). However, the documents or web pages retrieved by upstream IR systems are usually long. Therefore, it is difficult to concatenate all the retrieved documents and feed them into LLMs to generate answers. Though some approaches manage to solve these problems by aggregating the answers supported by each reference as the final answers, this strategy neglects the potential relations between retrieved passages. A more straightforward way is to directly compress the retrieved documents into short input tokens or even dense vectors [191, 192, 193, 194].
To compress the retrieved references, an intuitive idea is to extract the most useful $K$ sentences from the retrieved documents. LeanContext [191] applies this method and trains a small model by reinforcement learning (RL) to select the top $K$ similar sentences to the queries. The researchers also augment this strategy by using a free open-sourced text reduction method for the rest sentences as a supplement. Instead of using RL-based methods, RECOMP [192] directly uses the probability or the match ratio of the generated answers to the golden answers as signals to build training datasets and tune the compressor model. For example, the sentence corresponding to the highest generating probability is the positive one while others are negative ones. Furthermore, FILCO [193] applies the “hindsight” methods, which directly align the prior distribution (the predicted importance probability distribution of sentences without knowing the gold answer) to the posterior distribution (the same distribution of sentences within knowing the gold answer) to tune language models to select sentences.
However, these extractive methods may lose potential intent among all references. Therefore, abstractive methods are proposed to summarize retrieved documents into short but concise summaries for downstream generation. These methods [194, 192] usually distill the summarizing abilities of LLMs to small models. For example, TCRA [194] leverages GPT-3.5-turbo to build abstractive compression datasets for MT5 model.
6.4 Analysis
With the rapid development of the above reader approaches, many researchers have begun to analyze the characteristics of retrieval-augmented LLMs:
$\bullet$ Liu et al. [195] find that the position of the relevant/golden reference has significant influences on the final generation performance. The performance is always better when the relevant reference is at the beginning or the end, which indicates the necessity of introducing a ranking module to order the retrieved knowledge.
$\bullet$ Ren et al. [196] observe that by applying retrieval augmentation generation strategy, LLMs can have a better awareness of their knowledge boundaries.
$\bullet$ Liu et al. [197] analyze different strategies of integrating retrieval systems and LLMs such as concatenate (i.e., concatenating all references for answer generation) and post fusion (i.e., aggregating the answers corresponding to each reference). They also explore several ways of combining these two strategies.
$\bullet$ Aksitov et al. [198] demonstrate that there exists an attribution and fluency tradeoff for retrieval-augmented LLMs: with more received references, the attribution of generated answers increases while the fluency decreases.
$\bullet$ Mallen et al. [199] argue that always retrieving references to support LLMs to generate answers hurts the question-answering performance. The reason is that LLMs themselves may have adequate knowledge while answering questions about popular entities and the retrieved noisy passages may interfere and bias the answering process. To overcome this challenge, they devise a simple strategy that only retrieves references while the popularity of entities in the query is quite low. By this means, the efficacy and efficiency of retrieval-augmented generation both improve.
6.5 Applications
Recently, researchers [200, 201, 202, 203, 204, 205] have applied the retrieval-augmented generation strategy to areas such as clinical QA, medical QA, and financial QA to enhance LLMs with external knowledge and to develop domain-specific applications. For example, ATLANTIC [201] adapts Atlas to the scientific domain to derive a science QA system. Besides, some approaches [206] also apply techniques in federated learning such as multi-party computation to perform personal retrieval-augmented generation with privacy protection.
Furthermore, to better facilitate the deployment of these retrieval-augmented generation systems, some tools or frameworks are proposed [178, 207, 208]. For example, RETA-LLM [178] breaks down the whole complex generation task into several simple modules in the reader pipeline. These modules include a query rewriting module for refining query intents, a passage extraction module for aligning reference lengths with LLM limitations, and a fact verification module for confirming the absence of fabricated information in the generated answers.
@@ 6.6 Limitations
Several IR systems applying the retrieval-augmented generation strategy, such as New Bing and Langchain, have already entered commercial use. However, there are also some challenges in this novel retrieval-augmented content generation system. These include challenges such as effective query reformulation, optimal retrieval frequency, correct document comprehension, accurate passage extraction, and effective content summarization. It is crucial to address these challenges to effectively realize the potential of LLMs in this paradigm.
7 Search Agent
With the development of LLMs, IR systems are also facing new changes. Among them, developing LLMs as intelligent agents has attracted more and more attention. This conceptual shift aims to mimic human browsing patterns, thereby enhancing the capability of these models to handle complex retrieval tasks. Empowered by the advanced natural language understanding and generation capabilities of LLMs, these agents can autonomously search, interpret, and synthesize information from a wide range of sources.
One way to achieve this ability is to design a pipeline that combines a series of modules and assigns different roles to them. Such a pre-defined pipeline mimics users’ behaviors on the web by breaking it into several sub-tasks which are performed by different modules. However, this kind of static agent cannot deal with the complex nature of users’ behavior sequences on the web and may face challenges when interacting with real-world environments. An alternative solution is to allow LLMs to freely explore the web and make interactions themselves, namely letting the LLM itself decide what action it will take next based on the feedback from the environment (or humans). These agents have more flexibility and act more like human beings.
7.1 Static Agent
To mimic human search patterns, a straightforward approach is to design a static system to browse the web and synthesize information step by step [209, 210, 211, 212, 213, 214]. By breaking the information-seeking process into multiple subtasks, they design a pipeline that contains various LLM-based modules in advance and assigns different subtasks to them.
LaMDA [209] serves as an early work of the static agent. It consists of a family of Transformer-based neural language models specialized for dialog, with up to 137B parameters, pre-trained on 1.56T tokens from public dialogue data and web text. The study emphasizes the model’s development through a static pipeline, encompassing large-scale pre-training, followed by strategic fine-tuning stages aimed at enhancing three critical aspects: dialogue quality, safety, and groundedness. It can integrate external IR systems for factual grounding. This integration allows LaMDA to access and use external and authoritative sources when generating responses. SeeKeR [210] also incorporates the Internet search into its modular architecture for generating more factual responses. It performs three sequential tasks: generating a search query, generating knowledge from search results, and generating a final response. GopherCite [213] uses a search engine like Google Search to find relevant sources. It then synthesizes a response that includes verbatim quotes from these sources as evidence, aligning the Gopher’s output with verified information. WebAgent [212] develops a series of tasks, including instruction decomposition and planning, action programming, and HTML summarization. It can navigate the web, understand and synthesize information from multiple sources, and execute web-based tasks, effectively functioning as an advanced search and interaction agent. WebGLM [211] designs an LLM-augmented retriever, a bootstrapped generator, and a human preference-aware scorer. These components work together to provide accurate web-enhanced question-answering capabilities that are sensitive to human preferences. Shi et al. [214] focus on enhancing the relevance, responsibility, and trustworthiness of LLMs in web search applications via an intent-aware generator, an evidence-sensitive validator, and a multi-strategy supported optimizer.
7.2 Dynamic Agent
Instead of statically arranging LLMs in a pipeline, WebGPT [24] takes an alternate approach by training LLMs to use search engines automatically. This is achieved through the application of a reinforcement learning framework, within which a simulated environment is constructed for GPT-3 models. Specifically, the WebGPT model employs special tokens to execute actions such as querying, scrolling through rankings, and quoting references on search engines. This innovative approach allows the GPT-3 model to use search engines for text generation, enhancing the reliability and real-time capability of the generated texts. A following study [215] has extended this paradigm to the domain of Chinese question answering. Besides, some works develop important benchmarks for interactive web-based agents [216, 217, 218]. For example, WebShop [217] aims to provide a scalable, interactive web-based environment for language understanding and decision-making, focusing on the task of online shopping. ASH (Actor-Summarizer-Hierarchical) prompting [219] significantly enhances the ability of LLMs on WebShop benchmark. It first takes a raw observation from the environment and produces a new, more meaningful representation that aligns with the specific goal. Then, it dynamically predicts the next action based on the summarized observation and the interaction history.
7.3 Limitations
Though the aspect of static search agents has been thoroughly studied, the literature on dynamic search agents remains limited. Some agents may lack mechanisms for real-time fact-checking or verification against authoritative sources, leading to the potential dissemination of misinformation. Moreover, since LLMs are trained on data from the Internet, they may inadvertently perpetuate biases present in the training data. This can lead to biased or offensive outputs and may collect unethical content from the web. Finally, as LLMs process user queries, there are concerns regarding user privacy and data security, especially if sensitive or personal information is involved in the queries.
@@ 8.1 Query Rewriter
$\bullet$ Rewriting queries according to ranking performance. A typical paradigm of prompting-based methods is providing LLMs with several ground-truth rewriting cases (optional) and the task description of query rewriting. Despite LLMs being capable of identifying potential user intents of the query [220], they lack awareness of the resulting retrieval quality of the rewritten query. The absence of this connection can result in rewritten queries that seem correct yet produce unsatisfactory ranking results. Although some existing studies have used reinforcement learning to adjust the query rewriting process according to generation results [100], a substantial realm of research remains unexplored concerning the integration of ranking results.
@@
$\bullet$ Achieving personalized query rewriting. LLMs offer valuable contributions to personalized search through their capacity to analyze user-specific data. In terms of query rewriting, with the excellent language comprehension ability of LLMs, it is possible to leverage them to build user profiles based on users’ search histories (e.g., issued queries, click-through behaviors, and dwell time). This empowers the achievement of personalized query rewriting for enhanced IR and finally benefits personalized search or personalized recommendation.